A man is walking away from a lamppost with a light source 6m above the ground. The man is 2m tall. How long is the mans shadow when he is 10 m from the lightpost. Hint use similar triangles>
Answer by checkley77(12844)

2/x=6/(10+x) cross multiply.
6x=2(10+x)
6x=20+2x
6x-2x=20
4x=20
x=20/4
x=5 m. is the length of the man's shadow.